OpenStudy (anonymous):

During the month of February, Fabulous Feet Shoe Mart sold 48 pairs of red loafers. After an ad campaign to boost sales, they sold 60 pairs in March. Find the percent of increase in sales. A. 29% B. 19% C. 25% D. 15%

OpenStudy (triciaal):

increase = after ad - before ad = 60 -48 = 12 percent means per 100 12/48 = x /100 solve for x

OpenStudy (anonymous):

y 12 divided by 48 explan plz